OpenStudy (anonymous):

solve by substitution, y=-3x + 1 and x 2y=-3

jigglypuff314 (jigglypuff314):

is there supposed to be a sign in between the x and 2y?

OpenStudy (anonymous):

no, no sign

jigglypuff314 (jigglypuff314):

so I'm going to assume that it's just multiplying... knowing that y = -3x + 1 plug in that into x 2y = -3 so you get 2x(-3x + 1) = -3 then distribute and solve for x...

jigglypuff314 (jigglypuff314):

but truthfully, I'm getting weird answers :P ( (1/6 ± √(19)/6) , (1/2 ± √(19)/2) ) in other words, x = 1/6 ± √(19) /6 y = 1/2 ± √(19) /2

OpenStudy (anonymous):

that's what i got but i check the answer and it is (1,-2)

jigglypuff314 (jigglypuff314):

then you were missing a sign :P x + 2y = -3 works to give you those answers :)

OpenStudy (anonymous):

thanks i checked all my signs and i was missing a + sign , i appreciate your help